OpenStudy (anonymous):

5304=p+(0.04)(9) find the value of p

OpenStudy (anonymous):

lets see.

OpenStudy (anonymous):

do you know how to do the first step?

OpenStudy (anonymous):

0.04 * 9 = 9*4/100 =36/100 p = 5304 - 36/100 p = (530400-36)/100 p = 530364/100 0 = 5303.64